OpenStudy (anonymous):

Find a particular solution to the differential equation y′′+4y=cos^2(2x)

OpenStudy (anonymous):

turn it into its trig id: \[\cos^2 (2x) = \frac{ 1 + \cos4x }{ 2 }\] the form of the particulap solution should be: \[y_p = c_1 + c_2\sin4x + c_3 \cos4x\] plug y_p in the left side and solve for c1, c2 and c3 let me know if you have questions ^_^

OpenStudy (anonymous):

so i take the 2nd derivative of each particular solution and plug it in the left side

OpenStudy (anonymous):

oh no just the sum of the particular soliition

OpenStudy (anonymous):

yes and add it to 4 times yp and equate to the left side to solve for c_1 c_2 and c_3 protip: when you have cos or sin on the right side, you usually always need to have c1sinx + c2cosx. but since we are dealing with the 2nd derivative, you can completely omit sinx for this problem. yp is actually: c_1 + c_2cos(4x) since the c for sinx will be 0

OpenStudy (anonymous):

if we have y'' + y' + 4y on left then we would have to consider sinx.

OpenStudy (anonymous):

if you want i can solve it to show how it goes: i'll leave sin because you usually always need it, but c_2 will be zero so we don't need it there. \[y_p = c_1 + c_2 \sin(4x) + c_3 \cos(4x)\]\[y''_p = -16c_2 \sin(4x) - 16c_3\cos(4x)\] \[y'' + 4y = \frac{ 1 }{ 2 } + \frac{ \cos(4x) }{ 2 }\] \[-16c_2 \sin(4x) - 16 c_3 \cos(4x) + 4[c_1 + c_2\sin(4x) + _3\cos(4x)] = \frac{ 1 }{ 2 } + \frac{ \cos(4x) }{ 2 }\]

OpenStudy (anonymous):

\[4c_1 = \frac{ 1 }{ 2 }\]\[-12c_2\sin(4x) = 0*\sin(4x)\]\[-12 c_3 \cos(4x) = \frac{ \cos(4x) }{ 2 }\] c_1 = 1/8 c_2 = 0 c_3 = -1/24 \[y_p = \frac{ 1 }{ 8 } - \frac{ 1 }{ 24 } \cos(4x)\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

glad i could help ^_^ i love these actually